# 25mm25mm25mm26mm

### v2

@@ 4.1. Improved local law at the edge of the spectrum
|()|(**)$\prec$| ( ∗ ∗ ) | ≺ $\Big{(}\frac{\rho^{3/2}}{n\eta^{3/2}}\mathcal{E}_{b}+\frac{\rho^{3/2}}{\sqrt{n\eta}}\mathcal{E}_{b}+\frac{\rho^{2}}{\sqrt{n}\gamma}+\frac{\rho^{2}}{n\eta\gamma}\Big{)}{\mathbf{E}}\big{|}R^{xy}_{t}\big{|}^{2m-1}$
@@
Moreover, we have the analgue of (5.35), i.e., for $b^{\prime}=\min\big{\{}b+\theta,~{}1\big{\}}$ with $100<\theta\leq\epsilon/$ 10,
~$\frac{\widetilde{\mathcal{E}}_{b}}{\widetilde{\mathcal{E}}_{b^{\prime}}}=\Big{(}\frac{\gamma}{\eta}\Big{)}^{b^{\prime}-b}\lesssim\Big{(}\frac{1}{\rho^{2}\eta}\Big{)}^{\theta},\qquad\frac{1}{\eta\gamma\mathcal{E}_{b^{\prime}}}\lesssim 1.$ (5.46)
@@
For notational simplicity, we assume that $|\eta_{1}|\sim|\eta_{2}|\sim\eta$ and $\rho_{1}\sim\rho_{2}\sim\rho$. The proof for general $\eta_{i}$ and $\rho_{i}$ is essentially the same with $\eta$, $\rho$ replaced by $\eta_{*}$, $\rho^{*}$ in the following. We apply Itô’s formula to $|{R_{t}}|^{2m}$ and perform the cumulant expansions as in (5.1)-(5.27). By a direct computation, we obtain that
@@
where we also used (3.22) and (5.8) in the last step. In general for $s\geq 2$, each partial derivative $\partial^{s}R_{t}$ contains $s+2$ Green functions. The resulting terms have two forms: 1) a product of $s-1$ Green function entries $\prod_{i=1}^{s-1}(G)_{\bm{u}_{i},\bm{v}_{i}}$ and one factor $(G_{1}A_{1}G_{2}A_{2}G_{1})_{\bm{u}_{s}\bm{v}_{s}}$; 2) a product of $s-2$ Green function entries $\prod_{i=1}^{s-2}(G)_{\bm{u}_{i},\bm{v}_{i}}$, one factor $(G_{1}A_{1}G_{2})_{\bm{u}_{s-1}\bm{v}_{s-1}}$, and another factor $(G_{2}A_{2}G_{1})_{\bm{u}_{s}\bm{v}_{s}}$, with $\bm{u}_{i},\bm{v}_{i}~{}(1\leq i\leq s)$ chosen from $\{{\bm{e}}_{a},{\bm{e}}_{B}\}$. Note that both choices of ${\bm{e}}_{a}$ and ${\bm{e}}_{B}$ appear exactly $s$ times. For instance, we list below several examples of $\partial^{2}R_{t}$ and $\partial^{3}R_{t}$, i.e.,
@@
$\big{|}{\mathcal{T}}^{(3)}_{3}\big{|}\prec$ $\Big{(}\frac{\rho^{2}}{n^{3/2}\eta^{2}\sqrt{\gamma}}+\frac{\rho^{3/2}}{n\eta^{3/2}\gamma}+\frac{\rho}{n\gamma^{2}}\Big{)}{\mathbf{E}}\big{|}{R_{t}}\big{|}^{2m-1}$
@@
From (3.10) we know that $\rho_{i}\gtrsim\eta_{i}^{1/3}$ for $|z_{i}|\leq 1+n^{-1/2+\tau}$ and $\ell:=\min_{i=1}^{2}\{|\eta_{i}|\rho_{i}\}\gtrsim n^{-1+\epsilon}$ with $100<\tau<\epsilon/$ 10. This further implies that $\rho_{*}=\min_{i=1}^{2}\rho_{i}\gtrsim n^{-1/4-\epsilon/4}$. To prove Theorem 3.5, we will focus on the regime $|z_{1}-z_{2}|^{2}\gtrsim\rho_{1}\eta_{1}+\rho_{2}\eta_{2}$, which also implies that $\gamma\gtrsim\frac{\eta_{*}}{\rho^{*}}$ from (3.23). For the complementary regime $|z_{1}-z_{2}|^{2}\lesssim\rho_{1}\eta_{1}+\rho_{2}\eta_{2}$, the desired local law follows directly from Proposition 5.1 without gaining from the $|z_{1}-z_{2}|$-decay.
@@
$\widehat{R_{t}}:=\langle\big{(}\mathrm{Im}\,G_{1}A_{1}\mathrm{Im}\,G_{2}-\widehat{M}^{A}_{12}\big{)}A_{2}\rangle,\qquad G_{i}:=G^{z_{i}}_{t}(\mathrm{i}\eta_{i})=(G^{z_{i}}_{t}-\mathrm{i}\eta_{i})^{-1},\quad i=1,2,$ (5.74)
@@
for any $|z_{i}|\leq 1+\tau$, $n\ell\gtrsim n^{\epsilon}$ and $|\rho_{i}|\gtrsim n^{-c_{*}}$ with $40\leq c_{*}\leq 1/4-\epsilon/$ 4.
@@
recalling that $\rho_{*}=n^{-c_{*}}$ with $40\leq c_{*}\leq 1/4-\epsilon/$ 4. Clearly $a_{k}$ is increasing with $a_{k}-a_{k-1}\geq\epsilon/2$. Replacing Proposition 4.4 with Proposition 4.8, the $k$-th iteration step is similar to the $k$-th step explained in the proof of Proposition 5.2. Using the new scales defined in (5.77), we obtain the analogue of (5.25), i.e.,
@@
where we used that $40\leq c_{*}\leq 1/4-\epsilon/$ 4. We stop the iterations when $\eta^{(k)}_{*}$ reaches $\eta_{*}=\min_{i=1}^{2}|\eta_{i}|\gtrsim n^{-1+\epsilon}$. In the worst case when $z_{1}$ is in the bulk and $z_{2}$ is at the edge with $c_{*}=n^{-1/4-\epsilon/4}$ and $\eta_{*}=n^{-1+\epsilon}$, the number of iterations needed is at most $O(1/\epsilon)$. This hence proves Theorem 3.5 modulo the proof of (5.75) that we will do below. ∎
@@
Using that $\rho_{*}\gtrsim n^{-1/4}$, it is straightforward to check that the above finer bounds (5.4) and (5.4) are good enough to prove (5.75), for terms with $5\leq k\leq 7$, $l\leq 2k-8$, as well as the third and fourth order terms: $\widehat{\mathcal{T}}^{(1,1,1)}_{3}$, $\widehat{\mathcal{T}}^{(1,2)}_{3}$, $\widehat{\mathcal{T}}^{(1,1,1,1)}_{4}$,$\widehat{\mathcal{T}}^{(1,1,2)}_{4}$, $\widehat{\mathcal{T}}^{(1,3)}_{4}$. It then suffices to check the remaining terms, i.e., $\widehat{\mathcal{T}}^{(3)}_{3}$, $\widehat{\mathcal{T}}^{(4)}_{4}$, $\widehat{\mathcal{T}}^{(2,2)}_{4}$.
@@
$\big{|}\widehat{\mathcal{T}}^{(3)}_{3}\big{|}\prec$ $\Big{(}\frac{\rho^{*}}{n^{3/2}(\eta_{*})^{2}\gamma}+\frac{\rho^{*}}{n\eta_{*}\gamma}+\frac{\sqrt{\rho^{*}}}{n^{2}(\eta_{*})^{5/2}\gamma}+\frac{1}{(n\gamma)^{3/2}\eta_{*}}\Big{)}{\mathbf{E}}\big{|}\widehat{R_{t}}\big{|}^{2m-1}$
@@ Appendix A Proof of several lemmas and estimates in Section 4
Fix any $|z_{i}|<10$ and $|\eta_{i}|<10~{}(i=1,2)$. On the imaginary axis, both the product $\beta_{+}\beta_{-}$ and sum $\beta_{+}+\beta_{-}$ are positive and
@@
$=$ ±$\pm\frac{m_{1}^{2}m_{2}^{2}\pm m_{1}m_{2}+\mathrm{Re}\,[z_{1}\overline{z_{2}}]u_{1}u_{2}-|z_{1}|^{2}|z_{2}|^{2}u_{1}^{2}u_{2}^{2}}{\beta_{+}\beta_{-}};$ (A.21)
@@
Decomposing the matrix $A$ as $A=\langle AE_{+}\rangle E_{+}+\langle AE_{-}\rangle E_{-}+A_{c}$, we obtain that

### v3

@@ 4.1. Improved local law at the edge of the spectrum
$|(**)|\prec$ $\Big(\frac{\rho^{3/2}}{n\eta^{3/2}}\mathcal{E}_{b}+\frac{\rho^{3/2}}{\sqrt{n\eta}}\mathcal{E}_{b}+\frac{\rho^{2}}{\sqrt{n}\gamma}+\frac{\rho^{2}}{n\eta\gamma}\Big){\mathbf{E}}\big|R^{xy}_{t}\big|^{2m-1}$
@@
Moreover, we have the analgue of (5.35), i.e., for $b^{\prime}=\min\big\{b+\theta,~1\big\}$ with $0<\theta\leq\epsilon/10$,
$\frac{\widetilde{\mathcal{E}}_{b}}{\widetilde{\mathcal{E}}_{b^{\prime}}}=\Big(\frac{\gamma}{\eta}\Big)^{b^{\prime}-b}\lesssim\Big(\frac{1}{\rho^{2}\eta}\Big)^{\theta},\qquad\frac{1}{\eta\gamma\mathcal{E}_{b^{\prime}}}\lesssim 1.$ (5.46)
@@
For notational simplicity, we assume that $|\eta_{1}|\sim|\eta_{2}|\sim\eta$ and $\rho_{1}\sim\rho_{2}\sim\rho$. The proof for general $\eta_{i}$ and $\rho_{i}$ is essentially the same with $\eta$, $\rho$ replaced by $\eta_{*}$, $\rho^{*}$ in the following. We apply Itô’s formula to $|{R_{t}}|^{2m}$ and perform the cumulant expansions as in (5.1)-(5.27). As explained below (5.27), for notational brevity, we will use $\partial$ to indicate either $\partial/\partial h_{aB}$ or $\partial/\partial h_{Ba}$, and with a slight abuse of notation $R_{t}$ indicates either the complex number $R_{t}$ or $\overline{R_{t}}$. By a direct computation, we obtain that
@@
where we also used (3.22) and (5.8) in the last step. In general for $s\geq 2$, each partial derivative $\partial^{s}R_{t}$ contains $s+2$ Green functions. The resulting terms have two forms: 1) a product of s1s-1 Green function entries $\prod_{i=1}^{s-1}(G)_{\bm{u}_{i},\bm{v}_{i}}$ and one factor $(G_{1}A_{1}G_{2}A_{2}G_{1})_{\bm{u}_{s}\bm{v}_{s}}$; 2) a product of s2s-2 Green function entries $\prod_{i=1}^{s-2}(G)_{\bm{u}_{i},\bm{v}_{i}}$, one factor $(G_{1}A_{1}G_{2})_{\bm{u}_{s-1}\bm{v}_{s-1}}$, and another factor $(G_{2}A_{2}G_{1})_{\bm{u}_{s}\bm{v}_{s}}$, with $\bm{u}_{i},\bm{v}_{i}~(1\leq i\leq s)$ chosen from $\{{\bm{e}}_{a},{\bm{e}}_{B}\}$. Note that both choices of ${\bm{e}}_{a}$ and ${\bm{e}}_{B}$ appear exactly $s$ times. For instance, we list below several examples of $\partial^{2}R_{t}$ and $\partial^{3}R_{t}$, i.e.,
@@
[…]
@@
From (3.10) we know that $\rho_{i}\gtrsim\eta_{i}^{1/3}$ for $|z_{i}|\leq 1+n^{-1/2+\tau}$ and $\ell:=\min_{i=1}^{2}\{|\eta_{i}|\rho_{i}\}\gtrsim n^{-1+\epsilon}$ with $0<\tau<\epsilon/10$. This further implies that $\rho_{*}=\min_{i=1}^{2}\rho_{i}\gtrsim n^{-1/4-\epsilon/4}$. To prove Theorem 3.5, we will focus on the regime $|z_{1}-z_{2}|^{2}\gtrsim\rho_{1}\eta_{1}+\rho_{2}\eta_{2}$, which also implies that $\gamma\gtrsim\frac{\eta_{*}}{\rho^{*}}$ from (3.23). For the complementary regime $|z_{1}-z_{2}|^{2}\lesssim\rho_{1}\eta_{1}+\rho_{2}\eta_{2}$, the desired local law follows directly from Proposition 5.1 without gaining from the $|z_{1}-z_{2}|$-decay.
@@
$\widehat{R_{t}}:=\langle\big(\mathrm{Im}\,G_{1}A_{1}\mathrm{Im}\,G_{2}-\widehat{M}^{A}_{12}\big)A_{2}\rangle,\qquad G_{i}=G^{z_{i}}_{t}(\mathrm{i}\eta_{i})=(G^{z_{i}}_{t}-\mathrm{i}\eta_{i})^{-1},\quad i=1,2,$ (5.74)
@@
for any $|z_{i}|\leq 1+\tau$, $n\ell\gtrsim n^{\epsilon}$, $\gamma\gtrsim\frac{\eta_{*}}{\rho^{*}}$, and $|\rho_{i}|\gtrsim n^{-c_{*}}$ with $0\leq c_{*}\leq 1/4-\epsilon/4$.
@@
recalling that $\rho_{*}=n^{-c_{*}}$ with $0\leq c_{*}\leq 1/4-\epsilon/4$. Clearly $a_{k}$ is increasing with $a_{k}-a_{k-1}\geq\epsilon/2$. Replacing Proposition 4.4 with Proposition 4.8, the $k$-th iteration step is similar to the $k$-th step explained in the proof of Proposition 5.2. Using the new scales defined in (5.77), we obtain the analogue of (5.25), i.e.,
@@
where we used that $0\leq c_{*}\leq 1/4-\epsilon/4$. We stop the iterations when $\eta^{(k)}_{*}$ reaches $\eta_{*}=\min_{i=1}^{2}|\eta_{i}|\gtrsim n^{-1+\epsilon}$. In the worst case when $z_{1}$ is in the bulk and $z_{2}$ is at the edge with $c_{*}=n^{-1/4-\epsilon/4}$ and $\eta_{*}=n^{-1+\epsilon}$, the number of iterations needed is at most $O(1/\epsilon)$. This hence proves Theorem 3.5 modulo the proof of (5.75) that we will do below. ∎
@@
Using that $\rho_{*}\gtrsim n^{-1/4}$, it is straightforward to check that the above finer bounds (5.4) and (5.4) are good enough to prove (5.75), for terms with $5\leq k\leq 7$, $l\leq 2k-8$, as well as the third and fourth order terms: $\widehat{\mathcal{T}}^{(1,1,1)}_{3}$, $\widehat{\mathcal{T}}^{(1,2)}_{3}$, $\widehat{\mathcal{T}}^{(1,1,1,1)}_{4}$$\widehat{\mathcal{T}}^{(1,1,2)}_{4}$, $\widehat{\mathcal{T}}^{(1,3)}_{4}$. It then suffices to check the remaining terms, i.e., $\widehat{\mathcal{T}}^{(3)}_{3}$, $\widehat{\mathcal{T}}^{(4)}_{4}$, $\widehat{\mathcal{T}}^{(2,2)}_{4}$.
@@
[…]
@@ Appendix A Proof of several lemmas and estimates in Section 4
Fix any $|z_{i}|<10$ and $0<|\eta_{i}|<10~(i=1,2)$. On the imaginary axis, both the product $\beta_{+}\beta_{-}$ and sum $\beta_{+}+\beta_{-}$ are positive and
@@
$=$ $\pm\frac{m_{1}^{2}m_{2}^{2}\pm m_{1}m_{2}+\mathrm{Re}\,[z_{1}\overline{z_{2}}]u_{1}u_{2}-|z_{1}|^{2}|z_{2}|^{2}u_{1}^{2}u_{2}^{2}}{\beta_{+}\beta_{-}};$ (A.21)
@@
Decomposing the matrix AA as $A=\langle AE_{+}\rangle E_{+}+\langle AE_{-}\rangle E_{-}+A_{c}$, we obtain that